Given the function defined in the table below, find the average rate of change, in simplest form, of the function over the interval 1\le x \le 5

1≤x≤5.



1
Expert's answer
2022-06-09T13:01:11-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & f(x) \\ \hline
 0 & 3 \\
 \hdashline
 1 & 9 \\
 \hdashline
 2 & 27 \\
 \hdashline
 3 & 81 \\
 \hdashline
 4 & 243 \\
 \hdashline
 5 & 729 \\
 \hdashline
\end{array}"

"\text{average rate of change}=\dfrac{f(5)-f(1)}{5-1}"

"=\dfrac{729-9}{5-1}=180"